In a volumetric analysis experiment, an acidic aqueous solution of methanol (CH3OH) is titrated with a solution of potassium dichromate (K2Cr2O7) according to the following balanced chemical equation:

2 K2Cr2O7(aq) + 8 H2SO4(aq) + 3 CH3OH(aq) →

2 Cr2(SO4)3(aq) + 11 H2O(ª) + 3 HCOOH(aq) + 2 K2SO4(aq)

What volume of 1.15 M K2Cr2O7 is required to titrate 1.49 g of CH3OH dissolved in 45.0 mL of solution?

a.

45 mL

b.

1940 mL

c.

60.6 mL

d.

26.9 mL

e.

863 mL
Answer #1

moles of methanol = mass / molar mass = (1.49/32) = 0.0465 ,

moles of K2Cr2O2 from the balanced equation = (2/3)x0.0465 = 0.031,

moles = Molarity x volume

0.031 = 1.15 x Vol (K2Cr2O7)

volume = 0.0269 liters or 26.9 mL [ 1L=1000 mL]

so correct option is D

PLEASE LEAVE THE POSITIVE FEEDBACK IF YOU LIKE THE SOLUTION